# A Survey on the Applications of Zero-Knowledge Proofs

### v1

@@ IV-F3 Applications
Aleo, a Layer 1 privacy blockchain mentioned in this paper above, boasts its own ZK identity verification protocol called zPass [91]. zPass is a straightforward implementation due to the nature of enshrined zero-knowledge proofs and private execution available in Aleo. However, zPass also enables users to obtain anonymous credentials from existing identity documents, facilitating real-world adoption without necessitating changes to the protocol. Like Galxe protocol, it allows for selective attribute disclosure, enabling users to prove identity assertions across multiple credentials.
IV-G Supply Chain/Enterprise Blockchain Privacy
Figure 12: Supply Chain features fulfilled by Blockchain and Zero-Knowledge Proofs
IV-G1 Motivation and Definition
@@
Supply chains and global enterprises are increasingly recognizing the transformative potential of blockchain technology for enhancing transparency, provenance, immutability, and accountability. However, this shift towards blockchain adoption in supply chain comes with significant privacy concerns. In a typical supply chain, stakeholders, including suppliers, manufacturers, distributors, and retailers, share sensitive data such as pricing, inventory levels, and production schedules. The public and immutable nature of open blockchains could expose proprietary or sensitive information to competitors or the public. This trade-off between the need for transparency and privacy requirements has led to the exploration of privacy-enhancing technologies within blockchain frameworks for supply chain and enterprise applications.
IV-G2 Methodology
@@
There are several strategic approaches for implementing privacy in supply chain and enterprise blockchains. Firstly, the adoption of permissioned blockchain architectures enables the creation of a controlled ecosystem where access is granted only to verified participants. This selective visibility is crucial for maintaining a competitive edge and for ensuring that relevant data is shared only among trusted stakeholders. Furthermore, the integration of ZKPs addresses the necessity for privacy by allowing participants to verify the authenticity and compliance of products without revealing detailed process histories or proprietary information. This technology is instrumental in convincing end consumers of product safety and quality, even in the absence of full process transparency. To tackle the intricate and global nature of modern supply chains, data segmentation and encryption techniques are employed to protect business-sensitive information. By only storing cryptographic hashes of the actual data on-chain and keeping the detailed records off-chain or encrypted, these methodologies significantly reduce the risk of exposing trade secrets and sensitive business strategies. Moreover, the use of smart contracts for privacy enforcement plays a pivotal role in automating compliance and access control based on pre-defined rules. This approach not only streamlines operations but also ensures that data disclosure is strictly governed by necessity and consent in the supply chain scenario. Therefore, incorporating these enhanced privacy measures can achieve the goal of protecting end consumer interests and maintaining mutual trust among supply chain participants, as depicted in Figure 12.
IV-G3 Applications
@@
qedit [92] provides privacy-enhancing technology for enterprise blockchains, enabling secure collaboration and data sharing among parties without revealing sensitive information. It utilizes ZKPs to ensure that transactions are valid while keeping the transaction content private. It’s designed for enterprises looking to monetize data assets safely, enhance business analytics, and derive actionable insights in a secure manner. The platform is cloud-hosted, highly scalable, and integrates easily with existing database systems for quick deployment. qedit features a configurable dashboard, advanced reporting, and real-time notifications to provide businesses with critical intelligence efficiently. It is aimed at transforming the way companies collaborate on sensitive data, ensuring privacy while enabling data-driven decisions.
@@
IV-H Proof of Reserves
IV-H1 Motivation and Definition
@@ IV-H3 Applications
Provisions [96] offers a groundbreaking solution to this challenge by employing ZK-PoR. This technology enables a Bitcoin exchange, or any cryptocurrency platform, to transparently demonstrate that it possesses sufficient funds to cover all its obligations to users without disclosing the exact amounts held or the identities of the account holders. Allowing an exchange to prove its liquidity addresses the common concern of potential insolvency. Furthermore, this approach minimizes the risk of sensitive financial data being exposed, which could be detrimental in terms of privacy breaches or malicious exploitation.
@@
V Non-Blockchain Applications
Use Case Projects
Proof of Identity Zero-Knowledge Proofs of Identity, zk-creds: Flexible Anonymous Credentials from zkSNARKs and Existing Identity Infrastructure
@@ V Non-Blockchain Applications
Other Applications Non-interactive Zero-Knowledge Arguments for Voting, PhotoProof: Cryptographic Image Authentication for Any Set of Permissible Transformations, Experimenting with Collaborative zk-SNARKs: Zero-Knowledge Proofs for Distributed Secrets, VeeDo, ZKP2P
TABLE VII: Overview of non-blockchain applications of Zero-Knowledge Proofs
V-A Proof of Identity
V-A1 Motivation and Definition
Although we have seen many blockchain-based proof of identity protocols, identity proving as a practical application of ZKP extends past the blockchain domain as well. Fundamentally, this methodology is derived from the ability to prove a statement about identity, specifically membership in a set or credential verification, without revealing any sensitive information about the identity. As we saw in the blockchain domain, the space of privacy-preserving verifiable credentials is a popular result of these properties. This also extends to non-decentralized technologies, as we will see in the following section. This application domain can be defined as Proof of Identity, utilizing the privacy properties of zero-knowledge verifiable computation.
V-A2 Methodology
These identity proving techniques utilize ZKP to validate a user’s identity or membership in a group without revealing any specific, sensitive details about the identity itself. Compared to other application domains, this methodology is far more pure in the sense that the only unifying factor among applications is the ZKP itself; a proof is generated by a prover on some identity credential and given to an identity verifier, who does not get any identity information other than the proof itself.
V-A3 Applications
One of the earliest papers on ZKP, first published in 1988, presents a novel identification scheme based on zero-knowledge proofs, providing a more efficient alternative to RSA-based schemes. [98]. In identification schemes, entity A proves their identity to entity B using some constant S in the form of a value or physical card, without enabling entity B to then falsify themselves as A afterwards. Traditional identification schemes utilize encryption and/or hashing along with credentials such as digital passwords, PINs, credit card chips, etc. This paper proposes a practical scheme where no sophisticated adversary is capable of cooperating with a dishonest verifier B to produce a falsified credential and pretend to be A. The methodology involves interactive proofs where an entity can demonstrate their identity by proving they know a secret key of their credentials without having to reveal the secret; simply submitting a proof of their knowledge of the secret is sufficient, with the secret acting as a digital signature unique to each individual. The paper outlines a directory-less scheme, meaning there is no need for a central repository of public keys or identities. The paper also proposes that such a scheme could be implemented in hypothetical “smart cards”, which could act as physical credentials that generate zero-knowledge identity proofs using microprocessors to facilitate everyday identity verification.
The zk-creds protocol leverages zero-knowledge proofs (specifically zk-SNARKs) to convert existing identity documents into anonymous credentials, thereby eliminating the need for credential issuers to hold signing keys [99]. This system contrasts with traditional methods where issuers sign credentials and identity documents for validation. By integrating with existing identity infrastructures like government-issued IDs or university diplomas, zk-creds transforms these traditional credentials into digital, anonymous, yet verifiable formats.
V-B Machine Learning
V-B1 Motivation and Definition
@@
In a world where AI-generated material increasingly mimics human-created information, the potential use of zero-knowledge cryptography could assist us in determining that a specific piece of content was produced by applying a specific model to a given input. If a zero-knowledge circuit representation is built for them, this could give a technique for verifying outputs from large language models like GPT-4 [100], text-to-image models, or any other models. The zero-knowledge quality of these proofs allows us to hide sections of the input or the model if necessary. A good example is enabling users to view the model’s inference results without knowing the details of the model and prove that this result really comes from a specific model and input.
Zero-knowledge machine learning (ZKML) is a means to protect data privacy during model training and inference. They enable a data owner or a model owner to demonstrate the accuracy of a computation (such as the prediction of a machine learning model) without exposing any information about the data or the computation itself. This is especially effective in circumstances involving sensitive data.
V-B2 Methodology
@@
V-B3 Applications
ZKPs are commonly employed in machine learning, notably privacy-preserving and federated learning. service providers (provers) can use ZKP to prove the correctness of their model predictions without revealing the model itself to prevent model theft. The current state of the art in zero-knowledge systems coupled with performant hardware is still a few orders of magnitude shy of proving something as massive as currently available large language models (LLMs), but there has been considerable progress in establishing proofs of smaller models.
In verifying machine learning model predictions, vCNN [101] uses commit-and-prove to combine the typical quadratic arithmetic program (QAP) with the polynomial QAP in pairing-based zero-knowledge proofs. It supports convolutional neural networks and validates them using polynomial multiplications.
However, by presenting a novel sumcheck technique, zkCNN [102] could prove fast Fourier transformations and convolutions in a linear prover time. It is verified that the convolutions directly using the sumcheck protocol and zkCNN are 34×faster than vCNN.
zkDL[103] is an innovative approach to meet the need for zero-knowledge proofs in deep learning training. The main challenge it addresses is verifying the nonlinear computations inherent in neural networks, especially the ReLU activation function and its backpropagation. By introducing zkReLU, zkDL can efficiently handle these non-arithmetic operations without resorting to polynomial approximations, which are usually computationally expensive and less accurate. The FAC4DNN utilized by zkDL is an arithmetic circuit that aggregates proofs from different layers and training steps. This design bypasses traditional sequential proof generation and greatly reduces computational and communication overheads. zkDL also achieves full compatibility with tensor structures and supports large-scale neural networks, with proofs generated in less than a second for each batch update for networks with up to 10 million parameters.
@@ V-B3 Applications
The research by Sanjam Garg et al.[104] explores the practical application of zero-knowledge proofs in verifying the training of machine learning models. It focuses on experimental settings to verify the feasibility of such proofs, taking into account computational complexity and scalability issues. Their work emphasizes the need to balance strong security guarantees with practical overhead, ensuring that the proofs are concise and the prover’s workload is manageable. By experimenting with various configurations and optimization techniques, it provides insights into making zero-knowledge proofs applicable to real-world deep learning applications, which is very suitable as a reference.
Another recent study, Kaizen[105], proposed a zero-knowledge proof system designed for deep neural networks (DNNs). It ensures that the submitted model is correctly trained on the submitted dataset without leaking any other information. Kaizen adopts a sum-check-based proof system optimized for the gradient descent algorithm and recursively combines these proofs in multiple iterations. This recursive combination ensures that the proof size and verifier time are independent of the number of iterations, making it highly scalable. Kaizen has the ability to handle large models such as VGG-11, and is more practical than general recursive proofs, significantly reducing prover runtime and memory overhead.
In the study of the reasoning process, zkLLM[106] focuses on providing zero-knowledge proofs for large language models, ensuring that the reasoning results are verifiable without revealing the underlying model or data. This is especially important for applications that require strong privacy protection, such as medical or financial fields where data confidentiality is critical. zkLLM uses advanced encryption technology to efficiently generate and verify proofs, and maintains the privacy of model parameters and input data through cryptographic means, enhancing trust in deployed AI systems.
@@
In addition to the issue of the trustworthiness of its predictions, the model’s reliance on opaque data sources has become a new challenge. People frequently desire to keep the inputs and parameters of machine learning models private and unknown to the general public. Because the input data may contain sensitive information such as personal finances or biometrics, and the model parameters may also hold critical secrets. To specify provers and verifiers, many ZKML tools represented by the ezkl library [107] can implement higher-level descriptions of machine learning models or other computational graph programs. A prover can demonstrate that a particular output is produced by running a specific neural network on a specific data set.
Modulus Labs [108] is also developing a zero-knowledge machine learning solution, enabling trustful integration of AI outputs into blockchain systems without revealing sensitive data or model details. This allows developers to retain ownership and control over their AI models rather than relying on centralized platforms to host and manage their models. And their new zero-knowledge prover, Remainder[109], is a fast AI prover for AI inference. It is based on a verifiable decision forest inference circuit using GKR protocol[110].
@@
Giza [111] is a machine learning platform built on StarkNet that can be used to deploy and extend machine learning models, as well as solve the interoperability issues faced in the use of cloud-based machine learning models, performance, and transparency issues. It uses the ONNX open format to improve interoperability. Through a series of common operators and file formats it defines, developers can freely use TensorFlow, PyTorch, Scikit-Learn, and other frameworks and tools. Since StarkNet runs on the Layer2 network, it can enable any decentralized application to achieve unlimited computing scale without affecting the composability and security of Ethereum. Therefore, you don’t have to worry about load and architecture issues when using Giza, and can concentrate on model development and iterate. Another benefit of being based on StarkNet is that most functions are managed by the blockchain, which makes it easier to monitor, track, and manage the model, greatly improving transparency.
V-C Other Applications
V-C1 Motivation and Definition
Zero-knowledge proofs (ZKPs) have a broad range of applications extending beyond blockchain technology. Their unique ability to verify the accuracy of information without revealing the data itself makes them valuable in various technological fields. This section further progresses into other diverse applications of ZKPs that may not necessarily fit into our predefined categorization. These applications leverage ZKPs to address challenges in other significant areas, including image authentication, decentralized voting systems, and secure multi-party processes by ensuring data privacy and computational integrity.
V-C2 Methodology
In particular, we focus on four different applications: 1) Image Authentication, 2) Secure Electronic Voting Systems, 3) Randomness Generation and Timelocks, and 4) Collaborative Computations. ZKPs can be used to authenticate images without revealing the original content while allowing certain transformations like cropping or rotation. This is achieved by attaching a zero-knowledge proof to each image, certifying its integrity through any permissible transformations. Secure voting systems can employ ZKPs to validate encrypted votes while simultaneously ensuring voter privacy and integrity. This allows the verification of the correctness of each vote while maintaining the confidentiality of the voter’s choice. In applications such as randomness generation and timelocks, ZKPs can provide proof of computational integrity and delay in a verifiable delay function (VDF). These systems combine a delay function with the zkSTARK protocol in order to create a VDF that is slow to compute but fast to verify. Other applications extend the use of zk-SNARKs to collaborative environments where multiple parties jointly produce a single proof over a distributed witness. It addresses the challenge of maintaining data privacy in multi-party computations. The approach involves adapting zk-SNARKs for multi-prover protocols using secret-sharing techniques. This allows the generation of a single proof that validates a computation over data distributed among several parties without revealing individual inputs.
V-C3 Applications
In the realm of digital media, journalism, and legal documentation, the authenticity of images is a critical concern. PhotoProof [112] addresses this by allowing images to be authenticated even after undergoing permissible transformations such as cropping or color adjustments. This application is particularly significant in legal scenarios where image evidence must remain untampered and in journalism, where the integrity of photographic evidence can be crucial in forming public opinion. PhotoProof utilizes ZKPs to ensure that any changes made to an image do not compromise its original authenticity, thereby maintaining the credibility of digital media.
Moreover, ZKPs can address the limitations of electronic voting systems, such as tampering or breaches in voter privacy. The integrity of the voting process in elections is of utmost importance in maintaining democratic principles. The proposed system for electronic voting using non-interactive zero-knowledge arguments [113] presents a secure method to validate votes while preserving voter anonymity. This application is significant in preventing electoral fraud and ensuring a fair and transparent voting process.
Generating unbiased randomness and implementing secure timelocks is essential in blockchain systems and cryptographic protocols. VeeDo [114] leverages ZKPs in VDFs to provide a reliable source of randomness that is resistant to manipulation, enhancing unpredictability and fairness in blockchain applications. In cryptographic protocols, VeeDo’s timelock feature can be utilized to secure information for a predetermined period by adding an extra layer of security to sensitive transactions or processes.
In fields like scientific research, business analytics, and healthcare, there is often a need to perform multi-party computations over shared data without revealing individual inputs. Collaborative zk-SNARKs [115] enables multiple parties to compute a joint result while ensuring that the privacy of individual institutions’ data is preserved.
For instance, hospitals and research institutions might collaborate on patient data for medical research while adhering to privacy laws set up by the government. Using collaborative zk-SNARKs, they can analyze aggregate data, such as treatment effectiveness or disease trends, without revealing individual patient details. This has profound implications for collaborative research, where data privacy is paramount, and for businesses that require secure multi-party computations for joint ventures or partnerships.
Next, ZKP2P [116] emerges as an innovative bridge that integrates traditional finance with decentralized finance through a trustless and privacy-centric framework. At its core, ZKP2P employs ZKPs to verify DomainKeys Identified Mail (DKIM) signatures in payment confirmation emails. This use of ZKPs is instrumental in ensuring the authenticity of transactions while maintaining the privacy of sensitive information. The architecture of ZKP2P is characterized by its incorporation of several key components: circuits for the secure verification of transaction details, smart contracts for managing trustless interactions, and ZK-Email Libraries essential for the generation of private proofs relating to email contents. The ZKP2P protocol is designed to be interoperable with prevalent Web2 payment systems, thereby bridging a significant gap between fiat and cryptocurrencies. By eliminating intermediaries and reducing transaction fees, ZKP2P aims to offer a more inclusive and efficient platform for crypto-fiat conversions with the potential to evolve into a global, on-chain, trustless payment network compatible with diverse financial applications in DeFi, NFTs, and gaming.
VI Conclusion and Future Work
In the rapidly evolving world of digital security and privacy, ZKPs have emerged as a revolutionary tool, offering a way to share proofs of computational integrity without revealing the computation’s input. This survey explored a wide range of ZKPs’ practical applications and use cases, showing their crucial role in advancing cryptographic solutions and enhancing privacy in digital interactions. By examining various use cases, from enhancing privacy in blockchain to securing verification processes, ZKPs have demonstrated their potential to meet some of the most pressing challenges in ensuring digital privacy and security. However, the journey from theoretical constructs to widely applied solutions for ZKPs is still ongoing. The survey uncovered the depth and breadth of ZKPs’ applicability, highlighting the need for further research, especially in optimizing their implementation and expanding their use cases.
In concluding our survey on the applications of ZKPs, we identify some promising directions for future research to broaden their practical use and applicability. Future works could explore lightweight ZKP protocols that are feasible for devices with limited computational capabilities, enabling secure, privacy-preserving communication in the IoT landscape. Furthermore, the integration of ZKPs with ML, particularly in the context of bigger and more complex models, presents space for breakthroughs. Investigating the ways in which ZKPs can facilitate privacy-preserving computation and verification of large machine learning models without exposing the underlying data or the model itself has the potential to transform data privacy in AI.
In the domain of Layer-2 blockchain scalability using ZKPs, future research includes improving SNARK proof generation times in order to enable universal synchronous composability among different Layer-2 rollups. Universal synchronous composability allows different rollups to access and update the same universal blockchain state by sharing a sequencer that posts state changes to Layer-1 – this would defragment the Layer-2 rollup landscape and consolidate blockchain liquidity and state data among all Layer-2 rollups. In order for this to be possible, hardware custom-built for a proving system will likely be required.
Future research could also include finding the implications of merging ZKPs into game theoretic mechanisms. Specifically, there could be more formal research on finding equilibria in privacy-preserving systems augmented by ZK. For example, placing truthful bids in auctions without revealing sensitive information about underlying assets, similar to the private DAO application seen in section IV-E3. Research could also be done on the feasability of financial price discovery in fully or partial private modes, for both auctions and exchanges. To extend from this, there is yet to be a ZK SNARK system which proves the valid execution of a private order-book exchange. This can also be extended to dark pools, which are privatized blockchain exchanges enabled by ZKPs, which can utilize the automated market maker (AMM) exchange model.
Another potential application of ZKPs, specifically within the blockchain space, could be the use of privacy-preserving proofs to mitigate the negative externalities of maximal extractable value (MEV). Some possibilities could include creating encrypted transaction mempools and relays using ZKPs.
Additional future applications and research could extend any of the application categories mentioned in this paper. There are set to be many more uses of ZKP for legal compliance in private financial applications, mitigation of AI risks such as deepfakes, and additional privacy enhancements in financial applications. As we advance our computational techniques, explore new applications, and deepen our theoretical understanding, ZKPs stand to significantly impact how privacy and security are achieved in the digital age.

### v2

@@ 4.6.3 Applications
Aleo, a Layer 1 privacy blockchain mentioned in this paper above, includes its own ZK identity verification protocol called zPass [5]. zPass is a straightforward implementation due to the nature of enshrined zero-knowledge proofs and private execution available in Aleo. However, zPass also enables users to obtain anonymous credentials from existing identity documents, facilitating real-world adoption without necessitating changes to the protocol. Like Galxe protocol, it allows for selective attribute disclosure, enabling users to prove identity assertions across multiple credentials.
4.7 Supply Chain/Enterprise Blockchain Privacy
4.7.1 Motivation and Definition
@@
Supply chains and global enterprises are increasingly adopting blockchain technology to improve transparency, provenance, immutability, and accountability. However, this shift towards blockchain adoption in supply chain comes with significant privacy concerns. In a typical supply chain, stakeholders, including suppliers, manufacturers, distributors, and retailers, share sensitive data such as pricing, inventory levels, and production schedules. The public and immutable nature of open blockchains could expose proprietary or sensitive information to competitors or the public. This trade-off between the need for transparency and privacy requirements has led to the exploration of privacy-enhancing technologies within blockchain frameworks for supply chain and enterprise applications.
4.7.2 Methodology
@@
Implementing privacy in supply chain and enterprise blockchains involves several strategic approaches. The use of permissioned blockchain architectures creates a controlled ecosystem where only verified participants can access data, ensuring that sensitive information is shared exclusively with trusted stakeholders. Additionally, zero-knowledge proofs (ZKPs) allow for the verification of product authenticity and compliance without disclosing detailed process histories, thereby maintaining consumer confidence in product safety and quality. To address the complex nature of modern supply chains, data segmentation and encryption techniques are employed. By storing only cryptographic hashes on-chain and keeping detailed records off-chain or encrypted, these methods protect business-sensitive information and reduce the risk of exposing trade secrets. Smart contracts further enhance privacy by automating compliance and access control based on predefined rules, ensuring that data disclosure is governed by necessity and consent. These measures collectively protect consumer interests and maintain trust among supply chain participants.
4.7.3 Applications
@@
qedit [84] provides privacy-enhancing technology for enterprise blockchains, enabling secure collaboration and data sharing among parties without revealing sensitive information. It utilizes ZKPs to ensure that transactions are valid while keeping the transaction content private. It is designed for enterprises looking to monetize data assets safely, enhance business analytics, and derive actionable insights in a secure manner. The platform is cloud-hosted, highly scalable, and integrates easily with existing database systems for quick deployment. qedit features a configurable dashboard, advanced reporting, and real-time notifications to provide businesses with critical intelligence efficiently. It is aimed at transforming the way companies collaborate on sensitive data, ensuring privacy while enabling data-driven decisions.
@@
Beyond provenance tracking, three additional enterprise privacy patterns have emerged. Private compliance proofs allow a regulated entity to demonstrate that a dataset satisfies a policy predicate–e.g., that all transactions fall below a reporting threshold–without revealing the underlying records [66]. B2B attestations enable one firm to prove aggregate properties of an invoice batch or shipment manifest (total value, item count, origin country) to a counterparty without disclosing line-item detail [67]. Selective disclosure credentials let an organization extract verifiable assertions from a signed credential bundle–for instance, proving ISO-9001 certification status without revealing the full audit report [86]. A shared bottleneck across these patterns is that circuit complexity scales with schema richness: the more fields a compliance predicate or credential format contains, the larger the constraint system and the longer the proving time.
Table 9: Enterprise ZKP patterns: statement proved, primary bottleneck, and deployment maturity.
Pattern Statement proved Bottleneck Maturity
Provenance tracking Product origin & chain of custody Hash-chain depth Pilot
Private compliance Policy predicate over private data Predicate circuit size Research
B2B attestation Aggregate properties of records Schema-dependent circuit Proof-of-concept
Selective disclosure Assertion from signed credential Signature verification in-circuit Pilot
4.8 Proof of Reserves
4.8.1 Motivation and Definition
@@ 4.8.3 Applications
Provisions [23] addresses this challenge using ZK-PoR. This technology enables a Bitcoin exchange, or any cryptocurrency platform, to transparently demonstrate that it possesses sufficient funds to cover all its obligations to users without disclosing the exact amounts held or the identities of the account holders. Allowing an exchange to prove its liquidity addresses the common concern of potential insolvency. This approach also reduces the risk of exposing sensitive financial data to privacy breaches or malicious exploitation.
@@
Despite its promise, ZK-PoR carries several under-examined assumptions and failure modes. First, completeness of liabilities is not guaranteed: a custodian can exclude accounts from the liability Merkle tree, understating obligations while the proof remains mathematically valid. Second, proofs are point-in-time snapshots–assets can be moved immediately after proof generation, so a passing proof does not imply ongoing solvency. Third, the address-to-reserve mapping is self-attested by the custodian; without an independent on-chain audit linking addresses to the proving entity, the proof may cover borrowed or temporarily parked assets. Finally, there exists an inherent auditability–privacy tradeoff: stronger privacy protections (e.g., hiding individual account balances) reduce the ability of external parties to detect account omissions, making it harder to catch the very fraud the mechanism is designed to prevent.
5 Non-Blockchain Applications
In this section, we begin with ML/AI as the main non-blockchain focus, then group the remaining application topics into a single category with a summary table.
Table 10: Overview of non-blockchain applications of Zero-Knowledge Proofs
Use Case Representative works/projects
@@ 5 Non-Blockchain Applications
Other applications Proofs of Identity, zk-creds; voting ZK arguments, PhotoProof, VeeDo, Collaborative zkSNARKS, ZKP2P
5.1 Machine Learning and AI
5.1.1 Motivation and Definition
@@
In a world where AI-generated material increasingly mimics human-created information, the potential use of zero-knowledge cryptography could assist us in determining that a specific piece of content was produced by applying a specific model to a given input. If a zero-knowledge circuit representation is built for them, this could give a technique for verifying outputs from large language models like GPT-4 [3], text-to-image models, or any other models. The zero-knowledge quality of these proofs allows us to hide sections of the input or the model if necessary. A good example is enabling users to view the model’s inference results without knowing the details of the model and prove that this result really comes from a specific model and input. Zero-knowledge machine learning (ZKML) is a means to protect data privacy during model training and inference. They enable a data owner or a model owner to demonstrate the accuracy of a computation (such as the prediction of a machine learning model) without exposing any information about the data or the computation itself. This is especially effective in circumstances involving sensitive data.
5.1.2 Methodology
@@
5.1.3 Applications
[…]
@@ 5.1.3 Applications
The research by Sanjam Garg et al.[32] explores the practical application of zero-knowledge proofs in verifying the training of machine learning models. It focuses on experimental settings to verify the feasibility of such proofs, taking into account computational complexity and scalability issues. Their work emphasizes the need to balance strong security guarantees with practical overhead, ensuring that the proofs are concise and the prover’s workload is manageable. By experimenting with various configurations and optimization techniques, it provides insights into making zero-knowledge proofs applicable to real-world deep learning applications, which is very suitable as a reference. Another recent study, Kaizen[2], proposed a zero-knowledge proof system designed for deep neural networks (DNNs). It ensures that the submitted model is correctly trained on the submitted dataset without leaking any other information. Kaizen adopts a sum-check-based proof system optimized for the gradient descent algorithm and recursively combines these proofs in multiple iterations. This recursive combination is designed so that proof size and verifier time do not grow linearly with the number of iterations, improving scalability. Kaizen has the ability to handle large models such as VGG-11, and is more practical than general recursive proofs, significantly reducing prover runtime and memory overhead.
@@
In the study of the reasoning process, zkLLM[99] focuses on providing zero-knowledge proofs for large language models, ensuring that the reasoning results are verifiable without revealing the underlying model or data. This is especially important for applications that require strong privacy protection, such as medical or financial fields where data confidentiality is critical. zkLLM uses advanced encryption technology to efficiently generate and verify proofs, and maintains the privacy of model parameters and input data through cryptographic means, enhancing trust in deployed AI systems. In addition to the issue of the trustworthiness of its predictions, the model’s reliance on opaque data sources has become a new challenge. People frequently desire to keep the inputs and parameters of machine learning models private and unknown to the general public. Because the input data may contain sensitive information such as personal finances or biometrics, and the model parameters may also hold critical secrets. To specify provers and verifiers, many ZKML tools represented by the ezkl library [109] can implement higher-level descriptions of machine learning models or other computational graph programs. A prover can demonstrate that a particular output is produced by running a specific neural network on a specific data set.
@@
Modulus Labs [62] is also developing a zero-knowledge machine learning solution, enabling trustful integration of AI outputs into blockchain systems without revealing sensitive data or model details. This allows developers to retain ownership and control over their AI models rather than relying on centralized platforms to host and manage their models. And their new zero-knowledge prover, Remainder[63], is a fast AI prover for AI inference. It is based on a verifiable decision forest inference circuit using GKR protocol[37]. Giza [35] is a machine learning platform built on StarkNet that can be used to deploy and extend machine learning models, as well as solve the interoperability issues faced in the use of cloud-based machine learning models, performance, and transparency issues. It uses the ONNX open format to improve interoperability. Through a series of common operators and file formats it defines, developers can freely use TensorFlow, PyTorch, Scikit-Learn, and other frameworks and tools. Since StarkNet runs on the Layer2 network, it can enable any decentralized application to achieve unlimited computing scale without affecting the composability and security of Ethereum. This design reduces some infrastructure burdens on developers, allowing greater focus on model development. Another benefit of being based on StarkNet is that most functions are managed by the blockchain, which makes it easier to monitor, track, and manage the model, greatly improving transparency.
5.2 Other Non-Blockchain Applications
Table 11: Summary of other non-blockchain applications of ZKPs
Area What is proved Representative works/projects
Identity and credentials Membership or attribute validity without disclosure Feige–Fiat–Shamir identification [29]; zk-creds [86]
Voting Ballot validity and tally correctness with privacy Non-interactive ZK arguments for voting [39]
Media authentication Valid transform of an image under allowed edits PhotoProof [68]
Randomness and timelocks Correctness of delay functions and derived randomness VeeDo [93]
Collaborative computation Correct joint computation over private inputs Collaborative zkSNARKS [72]
Web2–Web3 payment bridging Validity of off-chain payment evidence without revealing contents ZKP2P [110]
Proof of identity protocols use ZKPs to show possession of a credential, or membership in a set, without disclosing the credential itself. Early work formalized interactive identification based on proving knowledge of a secret [29]. More recent systems such as zk-creds turn existing identity documents into anonymous credentials that can be verified without issuers maintaining signing keys [86]. The remaining applications in Table 11 follow the same theme: prove that a rule was followed, without exposing the underlying data. PhotoProof targets image integrity under allowed edits [68]. Voting arguments target ballot correctness while keeping votes private [39]. VeeDo uses ZK machinery around delay functions for timelocks and randomness [93]. Collaborative zkSNARKS support joint analytics without sharing inputs [72], and ZKP2P proves properties of payment emails (e.g., DKIM signatures) to bridge payment rails and on-chain settlement without revealing email contents [110].
6 Conclusion and Future Work
Zero-knowledge proofs have matured from theoretical constructs into practical tools that bridge cryptographic research and real-world deployment. This survey examined their applications across blockchain privacy, scalable computation, identity, and emerging non-blockchain domains such as verifiable machine learning and regulatory compliance. Despite rapid progress, significant challenges remain in proof-system optimization, developer tooling, and broader adoption beyond the blockchain ecosystem. An especially promising intersection lies in merging ZKPs with game-theoretic mechanisms–for instance, enabling truthful auction bidding without revealing sensitive information about underlying assets, akin to the private DAO applications discussed in Section 4.5.3. Privacy-preserving financial price discovery, private order-book exchanges, dark pools using the AMM model, and encrypted transaction mempools for mitigating maximal extractable value (MEV) all represent open research frontiers.
We see several directions for future work in this area:
Standardization and interoperability: Efforts such as ZKProof.org have begun defining shared terminology and security guidelines, yet the ecosystem still lacks common intermediate representations and universal proof formats. Establishing interoperability standards across proof systems, together with reproducible benchmark suites, would lower integration costs and accelerate adoption by practitioners who currently face vendor lock-in.
Post-quantum considerations: Transparent-setup systems such as STARKs and FRI-based protocols already avoid pairing-based assumptions vulnerable to quantum adversaries. Concurrently, lattice- and hash-based SNARKs are under active investigation. A key open question is how to balance quantum resistance against proof size and on-chain verification cost, since post-quantum constructions currently incur substantial overhead.
Circuit security and verification tooling: Under-constrained circuits and hand-written bugs remain a significant practical risk, as discussed in the DSL landscape of Section 3.2. Formal verification of constraint systems, static-analysis tooling for ZK-specific DSLs, and automated under-constraint detection are critical for closing the gap between theoretical soundness guarantees and deployed system security.
Performance roadmap: Proof recursion and composition, folding schemes such as Nova and incremental verifiable computation (IVC), and hardware acceleration via FPGAs and ASICs (Section 3.4) together chart a path toward substantially lower-latency proving for latency-sensitive applications. Reducing prover costs by orders of magnitude would unlock use cases–such as universal synchronous composability among Layer-2 rollups–that are currently bottlenecked by proof generation time.